OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (3, –2) and is parallel to the line y = –4x + 1? x + 4y = 10 x – 4y = –10 4x – y = –10 4x + y = 10

OpenStudy (ajprincess):

The gradients of parallel lines are equal. So the gradient of the required line =-4. \(y-y_1=m(x-x_1)\) \(y-(-2)=-4(x-3)\) Can u do nw?

OpenStudy (mathlegend):

Also remember that when looking for a line that is parallel you are looking for the slope to be the same. So in the equation... y=mx+b m is the slope so it must be -4x

OpenStudy (mathlegend):

Just follow @ajprincess steps

OpenStudy (ajprincess):

Guess it is a typo. m=-4 and nt -4x. @MathLegend